# HOG-Diff: Higher-Order Guided Diffusion for Graph Generation

### v1

@@ 3 Higher-order Guided Diffusion Model
We now present our Higher-order Guided Diffusion (HOG-Diff) model, which enhances graph generation by exploiting higher-order structures. We begin by detailing a coarse-to-fine generation curriculum that incrementally constructs graphs, followed by the introduction of three essential supporting techniques: the diffusion bridge, spectral diffusion, and a denoising model, respectively. Finally, we provide theoretical evidence validating the efficacy of HOG-Diff.
Coarse-to-fine Generation
We draw inspiration from curriculum learning, a paradigm that mimics the human learning process by systematically organizing data in a progression from simple to complex (Soviany et al., 2022). Likely, an ideal graph generation curriculum should be a composition of multiple easy-to-learn and meaningful intermediate steps. Additionally, higher-order structures encapsulate rich structural properties beyond pairwise interactions that are crucial for various empirical systems (Huang et al., 2024). As a graph-friendly generation framework, HOG-Diff incorporates higher-order structures during the intermediate stages of forward diffusion and reverse generative processes, thereby realizing a coarse-to-fine generation curriculum.
To implement our coarse-to-fine generation curriculum, we introduce a key operation termed cell complex filtering (CCF). As illustrated in Fig. 2, CCF generates an intermediate state of a graph by pruning nodes and edges that do not belong to a given cell complex.
Definition 3.1 (Cell complex filtering).
@@
Given a graph $\bm{G}=(\bm{V},\bm{E})$ and its associated cell complex $\mathcal{S}$, the cell complex filtering operation produces a filtered graph $\bm{G}^{\prime}=(\bm{V}^{\prime},\bm{E}^{\prime})$ where $\bm{V}^{\prime}=\{v\in\bm{V}\mid\exists\;x_{\alpha}\in\mathcal{S}:v\in x_{\alpha}\}$, and $\bm{E}^{\prime}=\{(u,v)\in\bm{E}\mid\exists\;x_{\alpha}\in\mathcal{S}:u,v\in x_{\alpha}\}$.
This filtering operation is a pivotal step in decomposing the graph generation task into manageable sub-tasks, with the filtered results serving as natural intermediaries in hierarchical graph generation. The overall framework of our proposed framework is depicted in Fig. 1. Specifically, the forward and reverse diffusion processes in HOG-Diff are divided into $K$ hierarchical time windows, denoted as $\{[\tau_{k-1},\tau_{k}]\}_{k=1}^{K}$, where $0=\tau_{0}<\cdots<\tau_{k-1}<\tau_{k}<\cdots<\tau_{K}=T$. Our sequential generation progressively denoises the higher-order skeletons. First, we generate coarse-grained higher-order skeletons, and subsequently refine them into finer pairwise relationships, simplifying the task of capturing complex graph distributions. This coarse-to-fine approach inherently aligns with the hierarchical nature of many real-world graphs, enabling smoother training and improved sampling performance.
@@
Formally, our generation process factorizes the joint distribution of the final graph $\bm{G}_{0}$ into a product of conditional distributions across these time windows:
$p(\bm{G}_{0})=p(\bm{G}_{0}|\bm{G}_{\tau_{1}})p(\bm{G}_{\tau_{1}}|\bm{G}_{\tau_{2}})\cdots p(\bm{G}_{\tau_{K-1}}|\bm{G}_{T}).$ (5)
Here, intermediate states $\bm{G}_{\tau_{1}},\bm{G}_{\tau_{2}},\cdots,\bm{G}_{\tau_{K-1}}$ represents different levels of cell filtering results of the corresponding graph. Our approach aligns intermediate stages of the diffusion process with realistic graph representations, and the ordering reflects a coarse-to-fine generation process.
The forward process introduces noise in a stepwise manner while preserving intermediate structural information. During each time window $[\tau_{k-1},\tau_{k}]$, the evolution of the graph is governed by the following forward SDE:
@@
Reversing this process enables the model to generate authentic samples with desirable higher-order information. The reverse-time SDE corresponds to Eq. 6 is as follows:
$\begin{split}\mathrm{d}\bm{G}_{t}^{(k)}=&\left[\mathbf{f}_{k,t}(\bm{G}_{t}^{(k)})-g_{k,t}^{2}\nabla_{\bm{G}_{t}^{(k)}}\log p_{t}(\bm{G}_{t}^{(k)})\right]\mathrm{d}\bar{t}\\ &+g_{k,t}\mathrm{d}\bar{\bm{W}}_{t}.\end{split}$ (7)
Instead of using higher-order information as a direct condition, HOG-Diff employs it to guide the generation process through multiple steps. This strategy allows the model to incrementally build complex graph structures while maintaining meaningful structural integrity at each stage. Moreover, integrating higher-order structures into graph generative models improves interpretability by allowing analysis of their significance in shaping the graph’s properties.
Diffusion Bridge Process
We realize the guided diffusion based on the generalized Ornstein-Uhlenbeck (GOU) process (Ahmad, 1988; Luo et al., 2023b), a stationary Gaussian-Markov process characterized by its mean-reverting property. Over time, the marginal distribution of the GOU process stabilizes around a fixed mean and variance, making it well-suited for stochastic modelling with terminal constraints. The GOU process $\mathbb{Q}$ is governed by the following SDE:
$\mathbb{Q}:\mathrm{d}\bm{G}_{t}=\theta_{t}(\bm{\mu}-\bm{G}_{t})\mathrm{d}t+g_{t}(\bm{G}_{t})\mathrm{d}\bm{W}_{t},$ (8)
where $\bm{\mu}=\bm{G}_{\tau_{k}}$ is the target terminal state, $\theta_{t}$ denotes a scalar drift coefficient and $g_{t}$ represents the diffusion coefficient. To ensure the process remains analytically tractable, $\theta_{t}$ and $g_{t}$ are constrained by the relationship $g_{t}^{2}/\theta_{t}=2\sigma^{2}$ (Luo et al., 2023b), where $\sigma^{2}$ is a given constant scalar. Under these conditions, its transition probability admits a closed-form solution:
$\begin{split}p(&\bm{G}_{t}\mid\bm{G}_{s})=\mathcal{N}(\mathbf{m}_{s:t},v_{s:t}^{2}\bm{I})\\ &=\mathcal{N}\left(\bm{\mu}+\left(\bm{G}_{s}-\bm{\mu}\right)e^{-\bar{\theta}_{s:t}},\sigma^{2}(1-e^{-2\bar{\theta}_{s:t}})\bm{I}\right).\end{split}$ (9)
[…]
The Doob’s $h$-transform can modify an SDE such that it passes through a specified endpoint. When applied to the GOU process, this eliminates variance in the terminal state, driving the diffusion process toward a Dirac distribution centered at $\bm{G}_{\tau_{k}}$ (Heng et al., 2021; Yue et al., 2024).
Proposition 3.2.
Let $\bm{G}_{t}$ evolve according to the generalized OU process in Eq. 8, subject to the terminal conditional $\bm{\mu}=\bm{G}_{\tau_{k}}$. The conditional marginal distribution $p(\bm{G}_{t}\mid\bm{G}_{\tau_{k}})$ then evolves according to the following SDE:
$\mathrm{d}\bm{G}_{t}=\theta_{t}\left(1+\frac{2}{e^{2\bar{\theta}_{t:\tau_{k}}}-1}\right)(\bm{G}_{\tau_{k}}-\bm{G}_{t})\mathrm{d}t+g_{k,t}\mathrm{d}\bm{W}_{t}.$
The conditional transition probability $p(\bm{G}_{t}\mid\bm{G}_{\tau_{k-1}},\bm{G}_{\tau_{k}})$ has analytical form as follows:
@@
We can directly use the closed-form solution in Prop. 3.2 for one-step forward sampling without performing multi-step forward iteration using the SDE. The reverse-time dynamics of the conditioned process can be derived using the theory of SDEs and take the following form:
$\mathrm{d}\bm{G}_{t}=\left[\mathbf{f}_{k,t}(\bm{G}_{t})-g_{k,t}^{2}\nabla_{\bm{G}_{t}}\log p(\bm{G}_{t}|\bm{G}_{\tau_{k}})\right]\mathop{}\!\mathrm{d}\bar{t}+g_{k,t}\mathop{}\!\mathrm{d}\bar{\bm{W}}_{t},$
where $\mathbf{f}_{k,t}(\bm{G}_{t})=\theta_{t}\left(1+\frac{2}{e^{2\bar{\theta}_{t:\tau_{k}}}-1}\right)(\bm{G}_{\tau_{k}}-\bm{G}_{t})$.
Spectral Diffusion
Generating graph adjacency matrices presents several significant challenges. Firstly, the non-uniqueness of graph representations implies that a graph with $n$ vertices can be equivalently modelled by up to $n!$ distinct adjacency matrices. This ambiguity requires a generative model to assign probabilities uniformly across all equivalent adjacencies to accurately capture the graph’s inherent symmetry. Additionally, unlike densely distributed image data, graphs typically follow a Pareto distribution and exhibit sparsity (Ghavasieh & De Domenico, 2024), so that adjacency score functions lie on a low-dimensional manifold. Consequently, noise injected into out-of-support regions of the full adjacency space severely degrades the signal-to-noise ratio, impairing the training of the score-matching process. Even for densely connected graphs, isotropic noise distorts global message-passing patterns by encouraging message-passing on sparsely connected regions. Moreover, the adjacency matrix scales quadratically with the number of nodes, making the direct generation of adjacency matrices computationally prohibitive for large-scale graphs.
@@
To address these challenges, inspired by Martinkus et al. (2022); Luo et al. (2023a), we introduce noise in the eigenvalue domain of the graph Laplacian matrix $\bm{L}=\bm{D}-\bm{A}$, instead of the adjacency matrix $\bm{A}$, where $\bm{D}$ denotes the diagonal degree matrix. As a symmetric positive semi-definite matrix, the graph Laplacian can be diagonalized as $\bm{L}=\bm{U}\bm{\Lambda}\bm{U}^{\top}$. Here, the orthogonal matrix $\bm{U}=[\bm{u}_{1},\cdots,\bm{u}_{n}]$ comprises the eigenvectors, and the diagonal matrix $\bm{\Lambda}=\operatorname{diag}(\lambda_{1},\cdots,\lambda_{n})$ holds the corresponding eigenvalues. The relationship between the Laplacian spectrum and the graph’s topology has been extensively explored (Chung, 1997). For instance, the low-frequency components of the spectrum capture the global structural properties such as connectivity and clustering, whereas the high-frequency components are crucial for reconstructing local connectivity patterns. Therefore, the target graph distribution $p(\bm{G}_{0})$ represents a joint distribution of $\bm{X}_{0}$ and $\bm{\Lambda}_{0}$, exploiting the permutation invariance and structural robustness of the Laplacian spectrum. Consequently, we split the reverse-time SDE into two parts that share drift and diffusion coefficients as
[…]
@@
Here, the superscript of $\bm{X}^{(k)}_{t}$ and $\bm{\Lambda}^{(k)}_{t}$ are dropped for simplicity, and $\mathbf{f}_{k,t}$ is determined according to Prop. 3.2.
@@
To approximate the score functions $\nabla_{\bm{X}_{t}}\log p_{t}(\bm{G}_{t}|\bm{G}_{\tau_{k}})$ and $\nabla_{\bm{\Lambda}_{t}}\log p_{t}(\bm{G}_{t}|\bm{G}_{\tau_{k}})$, we employ a neural network $\bm{s}^{(k)}_{\bm{\theta}}(\bm{G}_{t},\bm{G}_{\tau_{k}},t)$, composed of a node ($\bm{s}^{(k)}_{\bm{\theta},\bm{X}}(\bm{G}_{t},\bm{G}_{\tau_{k}},t)$) and a spectrum ($\bm{s}^{(k)}_{\bm{\theta},\bm{\Lambda}}(\bm{G}_{t},\bm{G}_{\tau_{k}},t)$) output, respectively. The model is optimized by minimizing the loss function:
$\ell^{(k)}(\bm{\theta})=$ $\mathbb{E}_{t,\bm{G}_{t},\bm{G}_{\tau_{k-1}},\bm{G}_{\tau_{k}}}\{\omega(t)[c_{1}\|\bm{s}^{(k)}_{\bm{\theta},\bm{X}}-\nabla_{\bm{X}}\log p_{t}(\bm{G}_{t}|\bm{G}_{\tau_{k}})\|_{2}^{2}$
$+$ $c_{2}||\bm{s}^{(k)}_{\bm{\theta},\bm{\Lambda}}-\nabla_{\bm{\Lambda}}\log p_{t}(\bm{G}_{t}|\bm{G}_{\tau_{k}})||_{2}^{2}]\},$ (11)
@@
where $\omega(t)$ is a positive weighting function, and $c_{1},c_{2}$ controls the relative importance of vertices and spectrum. The training procedure is detailed in Alg. 1 in the Appendix.
We sample $(\bm{X}_{\tau_{K}},\bm{\Lambda}_{\tau_{K}})$ from the prior distribution and uniformly sample $\bm{U}_{0}$ from the observed eigenvector matrices. The generation process involves multi-step diffusion to produce samples $(\hat{\bm{X}}_{\tau_{K-1}},\hat{\bm{\Lambda}}_{\tau_{K-1}}),\cdots,(\hat{\bm{X}}_{\tau_{1}},\hat{\bm{\Lambda}}_{\tau_{1}}),(\hat{\bm{X}}_{0},\hat{\bm{\Lambda}}_{0})$ in sequence by reversing the diffusion bridge, where the endpoint of one generation step serves as the starting point for the next. Finally, plausible samples with higher-order structures $\hat{\bm{G}}_{0}=(\hat{\bm{X}}_{0},\hat{\bm{L}}_{0}=\bm{U}_{0}\hat{\bm{\Lambda}}_{0}\bm{U}_{0}^{\top})$ can be reconstructed. The complete sampling procedure is outlined in Alg. 2 within the Appendix.
Table 1: Comparison of different methods based on molecular datasets. The best results for the first three metrics are highlighted in bold.
Methods QM9 ZINC250k
[…]
GraphAF 0.020 5.268 67.00 100.00 94.51 88.83 0.044 16.289 68.00 100.00 99.10 100.00
GraphAF+FC 0.021 5.625 74.43 100.00 86.59 89.57 0.044 16.023 68.47 100.00 98.64 99.99
GraphDF 0.063 10.816 82.67 100.00 97.62 98.10 0.176 34.202 89.03 100.00 99.16 99.99
GraphDF+FC 0.064 10.928 93.88 100.00 98.32 98.54 0.177 33.546 90.61 100.00 99.63 100.00
MoFlow 0.017 4.467 91.36 100.00 98.65 94.72 0.046 20.931 63.11 100.00 99.99 99.99
EDP-GNN 0.005 2.680 47.52 100.00 99.25 86.58 0.049 16.737 82.97 100.00 99.79 99.99
GraphEBM 0.003 6.143 8.22 100.00 99.25 85.48 0.212 35.471 5.29 99.96 98.79 99.99
GDSS 0.003 2.900 95.72 100.00 98.46 86.27 0.019 14.656 97.01 100.00 99.64 100.00
DiGress 0.0005 0.360 99.00 100.00 96.66 33.40 0.082 23.060 91.02 100.00 81.23 100.00
HOG-Diff 0.0003 0.172 98.74 100.00 97.01 75.12 0.001 1.533 98.56 100.00 99.96 99.43
Denoising Network Architecture
@@
We design a neural network $\bm{s}^{(k)}_{\bm{\theta}}(\bm{G}_{t},\bm{G}_{\tau_{k}},t)$ to estimate score functions in Sec. 3. Standard graph neural networks designed for classical tasks such as graph classification and link prediction may be inappropriate for graph distribution learning due to the immediate real-number graph states and the complicated requirements. For example, an effective model for molecular graph generation should capture local node-edge dependence for chemical valency rules and attempt to recover global graph patterns like edge sparsity, frequent ring subgraphs, and atom-type distribution.
@@
To achieve this, we introduce a unified denoising network that explicitly integrates node and spectral representations. As illustrated in Fig. 6 of Appendix, the network comprises two different graph processing modules: a standard graph convolution network (GCN) (Kipf & Welling, 2017) for local feature aggregation and a graph transformer network (ATTN)  (Dwivedi & Bresson, 2021; Vignac et al., 2023) for global information extraction. The outputs of these modules are fused with time information through a Feature-wise Linear Modulation (FiLM) layer (Perez et al., 2018). The resulting representations are concatenated to form a unified hidden embedding. This hidden embedding is further processed through separate multilayer perceptrons (MLPs) to produce predictions for $\nabla_{\bm{X}}\log p(\bm{G}_{t}|\bm{G}_{\tau_{k}})$ and $\nabla_{\bm{\Lambda}}\log p(\bm{G}_{t}|\bm{G}_{\tau_{k}})$, respectively. It is worth noting that our graph noise prediction model is permutation equivalent as each component of our model avoids any node ordering-dependent operations. Our model is detailed in Sec. D.1.
Theoretical Analysis

### v2

@@ 3 Higher-order Guided Diffusion Model
We now present our Higher-order Guided Diffusion (HOG-Diff) model, which enhances graph generation by exploiting higher-order structures. We first describe our coarse-to-fine generation framework, followed by an introduction to the supporting diffusion bridge technique. Finally, we provide theoretical evidence to validate the efficacy of HOG-Diff.
3.1 Coarse-to-fine Framework with Topological Filtering
We draw inspiration from curriculum learning, a paradigm that mimics human learning by systematically organizing data in a progression from simple (coarse) to complex (fine) (Abbe et al., 2021; Soviany et al., 2022). Specifically, we model coarse intermediary structures as higher-order cells, which encapsulate rich structural properties beyond pairwise interactions (Battiston et al., 2020). These cells can be obtained by lifting the original graph and retaining associated 2-faces as the higher-order skeleton. Our generative processes then follow a curriculum to progressively generate graphs, starting with higher-order cells and gradually refining them into the full complex graph.
Figure 2: Cell Complex transformations. (a) An example graph. (b) Lifting: closed 2D disks are glued to the boundary of the rings to form the 2-cell complex. (c) The resulting cell complex and the corresponding homeomorphisms to the closed balls for three representative cells of different dimensions in the complex. (d) Black elements represent higher-order structures extracted through 2-cell filtering, while grey elements denote corresponding peripheral structures pruned by the filtering operation.
To implement our coarse-to-fine generation framework, we first introduce a key operation termed cell complex filtering (CCF). The filtering operation decomposes the graph generation task into hierarchically structured and manageable subtasks.
Proposition 2 (Cell complex filtering).
@@ 3.1 Coarse-to-fine Framework with Topological Filtering
Given a graph $\bm{G}=(\bm{V},\bm{E})$ and its associated cell complex $\mathcal{S}=\cup_{\alpha}x_{\alpha}$ (obtained via lifting). The $p$-cell complex filtering operation defines a filtered graph $\bm{G}_{p}=(\bm{V}_{p},\bm{E}_{p})$, where $\bm{V}_{p}=\{v\in\bm{V}\mid\exists\;x_{\alpha}\text{with}\dim(x_{\alpha})=p:v\in\overline{x_{\alpha}}\}$, and $\bm{E}_{p}=\{(u,v)\in\bm{E}\mid\exists\;x_{\alpha}\text{with}\dim(x_{\alpha})=p:\{u,v\}\subseteq\overline{x_{\alpha}}\}$. Here, $\overline{x_{\alpha}}$ denotes the closure of $x_{\alpha}$.
As illustrated in Fig. 2, we first lift the graph to a cell complex and then apply CCF to generate intermediate states by pruning nodes and edges that do not belong to higher-order cells. In practice, CCF offers a substantial speedup, as it avoids the expensive enumeration of all cells required by lifting, see App. D for details. Based on the filtering results, the diffusion process is structured into $K$ hierarchical time windows, denoted as $\{[\tau_{k-1},\tau_{k}]\}_{k=1}^{K}$, where $0=\tau_{0}<\cdots<\tau_{k-1}<\tau_{k}<\cdots<\tau_{K}=T$, with the filtered results serving as natural intermediaries in the hierarchical generation process.
@@
The overall framework of HOG-Diff is depicted in Fig. 1. In general, we first generate coarse-grained higher-order skeletons and subsequently refine them into finer pairwise relationships, thereby simplifying the task of capturing complex graph distributions. Formally, our generation process factorizes the joint distribution of the final graph $\bm{G}_{0}$ into a product of conditional distributions across these time windows:
$p(\bm{G}_{0})=p(\bm{G}_{0}|\bm{G}_{\tau_{1}})p(\bm{G}_{\tau_{1}}|\bm{G}_{\tau_{2}})\cdots p(\bm{G}_{\tau_{K-1}}|\bm{G}_{T}).$
Here, the intermediate states $\bm{G}_{\tau_{K-1}},\cdots,\bm{G}_{\tau_{2}},\bm{G}_{\tau_{1}}$ represent progressively finer cell complex filtered graph representations, aligning intermediate diffusion stages with realistic hierarchical graph structures. This coarse-to-fine approach enables our model to first focus on fundamental topological structures and then add finer connectivity, inherently aligning with the hierarchical nature of many empirical systems. Consequently, our model benefits from smoother training and improved sampling performance (see Sec. 3.3 for theoretical analysis).
To ensure smooth transitions between intermediate states within each interval $\tau_{k-1},\tau_{k}]$, the graph evolves according to the general form of a diffusion bridge process (see Sec. 3.2 for details):
@@
The forward diffusion process introduces noise in a stepwise manner while preserving intermediate structural information. Reversing this process enables the model to generate authentic samples with desirable higher-order information. Moreover, integrating higher-order structures into graph generative models improves interpretability by allowing analysis of their significance in shaping the graph’s properties. Rather than directly conditioning on higher-order information, HOG-Diff employs it incrementally as a guiding structure. This strategy allows the model to build complex graph structures progressively, while maintaining meaningful structural integrity at each stage.
3.2 Guided Generation via Diffusion Bridge Process
As a building block of our generative framework, we leverage a class of diffusion processes with fixed terminal states, namely the generalized Ornstein-Uhlenbeck (GOU) bridge, to realize the proposed guided diffusion process in Eq. 3, while enabling simulation-free training.
The generalized Ornstein-Uhlenbeck (GOU) process (Ahmad, 1988; Luo et al., 2023b), also known as the time-varying OU process, is a stationary, Gaussian-Markov process characterized by its mean-reverting property. Specifically, the marginal distribution of the GOU process asymptotically approaches a fixed mean and variance. The GOU process is governed by the following SDE:
$\mathrm{d}\bm{G}_{t}=\theta_{t}(\bm{\mu}-\bm{G}_{t})\mathrm{d}t+g_{t}\mathrm{d}\bm{W}_{t},$ (4)
where $\bm{\mu}$ is the target terminal state, $\theta_{t}$ denotes a scalar drift coefficient and $g_{t}$ represents the diffusion coefficient. For analytical tractability, $\theta_{t}$ and $g_{t}$ are constrained by $g_{t}^{2}/\theta_{t}=2\sigma^{2}$ (Luo et al., 2023b), where $\sigma^{2}$ is a fixed constant, yielding a closed-form transition probability:
$p(\bm{G}_{t}\mid\bm{G}_{s})=\mathcal{N}(\mathbf{m}_{s:t},v_{s:t}^{2}\bm{I})=\mathcal{N}\left(\bm{\mu}+\left(\bm{G}_{s}-\bm{\mu}\right)e^{-\bar{\theta}_{s:t}},\sigma^{2}(1-e^{-2\bar{\theta}_{s:t}})\bm{I}\right).$ (5)
Here, $\bar{\theta}_{s:t}=\int_{s}^{t}\theta_{z}dz$. For notional simplicity, $\bar{\theta}_{0:t}$ is replaced by $\bar{\theta}_{t}$ when $s=0$.
Diffusion Bridge
Applying Doob’s $h$-transform (Doob & Doob, 1984) to the GOU process under the terminal condition $\bm{\mu}=\bm{G}_{\tau_{k}}$, we can derive the GOU bridge process as follows (we provide the detailed derivation of the bridge process in Sec. A.1):
The conditional transition probability admits an analytical form $p(\bm{G}_{t}|\bm{G}_{\tau_{k-1}},\bm{G}_{\tau_{k}})=\mathcal{N}(\bar{\mathbf{m}}_{t},\bar{v}_{t}^{2}\bm{I})$:
$\bar{\mathbf{m}}_{t}=\bm{G}_{\tau_{k}}+(\bm{G}_{\tau_{k-1}}-\bm{G}_{\tau_{k}})e^{-\bar{\theta}_{\tau_{k-1}:t}}\frac{v_{t:\tau_{k}}^{2}}{v_{\tau_{k-1}:\tau_{k}}^{2}},\quad\bar{v}_{t}^{2}={v_{\tau_{k-1}:t}^{2}v_{t:\tau_{k}}^{2}}/{v_{\tau_{k-1}:\tau_{k}}^{2}}.$ (7)
@@ 3.2 Guided Generation via Diffusion Bridge Process
The GOU bridge process eliminates variance in the terminal state by directing the diffusion toward a Dirac distribution centered at $\bm{G}_{\tau_{k}}$, making it well-suited for stochastic modelling with terminal constraints (Heng et al., 2021; Yue et al., 2024). Moreover, we can directly use the closed-form solution for one-step forward sampling without expensive SDE simulation. Note that the Brownian bridge process used in previous works (Wu et al., 2022) is a special case of the GOU bridge process when $\theta_{t}\rightarrow 0$ (see Sec. A.1).
Training and Sampling
@@
Classical graph diffusion approaches typically inject isotropic Gaussian noise directly into the adjacency matrices $\bm{A}$, leading to various fundamental challenges, such as permutation ambiguity, sparsity-induced signal degradation, and poor scalability (see Sec. C.2 for detailed discussion). To address these challenges, inspired by Luo et al. (2023a), we introduce noise in the spectra of the graph Laplacian $\bm{L}=\bm{D}-\bm{A}$, instead of the adjacency matrix $\bm{A}$, where $\bm{D}$ denotes the diagonal degree matrix. As a symmetric positive semi-definite matrix, the graph Laplacian can be diagonalized as $\bm{L}=\bm{U}\bm{\Lambda}\bm{U}^{\top}$. Here, the orthogonal matrix $\bm{U}=[\bm{u}_{1},\cdots,\bm{u}_{n}]$ comprises the eigenvectors, and the diagonal matrix $\bm{\Lambda}=\operatorname{diag}(\lambda_{1},\cdots,\lambda_{n})$ holds the corresponding eigenvalues. Therefore, the target graph distribution $p(\bm{G}_{0})$ represents a joint distribution of $\bm{X}_{0}$ and $\bm{\Lambda}_{0}$, exploiting the permutation invariance and structural robustness of the Laplacian spectrum.
Consequently, the GOU bridge process in Eq. 6, along with its time-reversed counterpart, can be formulated as the following system of SDEs for graph $\bm{G}$:
$\left\{\begin{aligned} \mathrm{d}\bm{X}_{t}=&\mathbf{f}_{k,t}(\bm{X}_{t})\mathrm{d}t+g_{k,t}\mathrm{d}\bar{\bm{W}}_{t}^{1}\\ \mathrm{d}\bm{\Lambda}_{t}=&\mathbf{f}_{k,t}(\bm{\Lambda}_{t})\mathrm{d}t+g_{k,t}\mathrm{d}\bar{\bm{W}}_{t}^{2}\end{aligned}\right.,\left\{\begin{aligned} \mathrm{d}\bm{X}_{t}=&\left[\mathbf{f}_{k,t}(\bm{X}_{t})-g_{k,t}^{2}\nabla_{\bm{X}}\log p_{t}(\bm{G}_{t}|\bm{G}_{\tau_{k}})\right]\mathrm{d}\bar{t}+g_{k,t}\mathrm{d}\bar{\bm{W}}_{t}^{1}\\ \mathrm{d}\bm{\Lambda}_{t}=&\left[\mathbf{f}_{k,t}(\bm{\Lambda}_{t})-g_{k,t}^{2}\nabla_{\bm{\Lambda}}\log p_{t}(\bm{G}_{t}|\bm{G}_{\tau_{k}})\right]\mathrm{d}\bar{t}+g_{k,t}\mathrm{d}\bar{\bm{W}}_{t}^{2}\end{aligned}\right.$.
@@
Here, the reverse-time dynamics of the bridge process are derived using the theory of SDEs, the superscript of $\bm{X}^{(k)}_{t}$ and $\bm{\Lambda}^{(k)}_{t}$ are dropped for simplicity, and $\mathbf{f}_{k,t}$ is determined according to Eq. 6.
@@
To approximate the score functions $\nabla_{\bm{X}_{t}}\log p_{t}(\bm{G}_{t}|\bm{G}_{\tau_{k}})$ and $\nabla_{\bm{\Lambda}_{t}}\log p_{t}(\bm{G}_{t}|\bm{G}_{\tau_{k}})$, we employ a neural network $\bm{s}^{(k)}_{\bm{\theta}}(\bm{G}_{t},\bm{G}_{\tau_{k}},t)$, which outputs predictions for both node-level ($\bm{s}^{(k)}_{\bm{\theta},\bm{X}}(\bm{G}_{t},\bm{G}_{\tau_{k}},t)$) and spectrum ($\bm{s}^{(k)}_{\bm{\theta},\bm{\Lambda}}(\bm{G}_{t},\bm{G}_{\tau_{k}},t)$) components. The network is optimized by minimizing:
$\ell^{(k)}(\bm{\theta})=\mathbb{E}_{t,\bm{G}_{t},\bm{G}_{\tau_{k-1}},\bm{G}_{\tau_{k}}}\{\omega(t)[c_{1}\|\bm{s}^{(k)}_{\bm{\theta},\bm{X}}-\nabla_{\bm{X}}\log p_{t}(\bm{G}_{t}|\bm{G}_{\tau_{k}})\|_{2}^{2}+c_{2}||\bm{s}^{(k)}_{\bm{\theta},\bm{\Lambda}}-\nabla_{\bm{\Lambda}}\log p_{t}(\bm{G}_{t}|\bm{G}_{\tau_{k}})||_{2}^{2}]\},$ (8)
@@
where $\omega(t)$ is a positive weighting function, and $c_{1},c_{2}$ controls the relative importance of vertices and spectrum. The training procedure is detailed in Alg. 1 in App. C.
In the inference procedure, we sample $(\hat{\bm{X}}_{\tau_{K}},\hat{\bm{\Lambda}}_{\tau_{K}})$ from the prior distribution and select $\hat{\bm{U}}_{0}$ as an eigenbasis drawn from the training set. Reverse diffusion is then applied across multiple stages to sequentially generate $(\hat{\bm{X}}_{\tau_{K-1}},\hat{\bm{\Lambda}}_{\tau_{K-1}}),\cdots,(\hat{\bm{X}}_{\tau_{1}},\hat{\bm{\Lambda}}_{\tau_{1}}),(\hat{\bm{X}}_{0},\hat{\bm{\Lambda}}_{0})$, where each stage is implemented via the diffusion bridge and initialized from the output of the previous step. Finally, plausible samples with higher-order structures can be reconstructed as $\hat{\bm{G}}_{0}=(\hat{\bm{X}}_{0},\hat{\bm{L}}_{0}=\hat{\bm{U}}_{0}\hat{\bm{\Lambda}}_{0}\hat{\bm{U}}_{0}^{\top})$. Further details of the spectral diffusion process and the complete sampling procedure are provided in App. C, while ablation studies comparing diffusion in the spectral domain versus the adjacency matrix are presented in Sec. F.1.
Score Network Architecture
@@
The score network plays a critical role in estimating the score functions required to reverse the diffusion process. Standard graph neural networks designed for classical tasks such as graph classification and link prediction may be inappropriate for graph distribution learning due to the complicated requirements. For example, an effective model for molecular graph generation should capture local node-edge dependence for chemical valency rules and attempt to recover global graph patterns like edge sparsity, frequent ring subgraphs, and atom-type distribution.
@@
To achieve this, we introduce a unified score network that explicitly integrates node and spectral representations. As illustrated in Fig. 7 of Appendix, the network comprises two different graph processing modules: a standard graph convolution network (GCN) (Kipf & Welling, 2017) for local feature aggregation and a graph transformer network (ATTN) (Dwivedi & Bresson, 2021; Vignac et al., 2023) for global information extraction. The outputs of these modules are fused with time information through a Feature-wise Linear Modulation (FiLM) layer (Perez et al., 2018), and the resulting representations are concatenated to form a unified hidden embedding. This hidden embedding is further processed by separate multilayer perceptrons (MLPs) to produce predictions for $\nabla_{\bm{X}}\log p(\bm{G}_{t}|\bm{G}_{\tau_{k}})$ and $\nabla_{\bm{\Lambda}}\log p(\bm{G}_{t}|\bm{G}_{\tau_{k}})$, respectively. It is worth noting that our score network is permutation equivalent, as each component of our model avoids any node ordering-dependent operations. Our model is detailed in Sec. C.3.
3.3 Theoretical Analysis